consider the linear transformation defined by F ( x , y , z ) = ( y z , x 2 ) . find f ( 2 , 3 , 4 )


1
Expert's answer
2021-05-13T07:07:52-0400

"F ( x , y , z ) = ( y -z , x-2 )\\
 F(2,3,4)=(3-4,2-2)=(-1,0)"